OpenStudy (unknownrandom):

Find y' by implicit differentiation. e^(x/y)=6x−6y

OpenStudy (unknownrandom):

I am really struggling with these problems. Can somebody walk me through it?

OpenStudy (irishboy123):

\[ \Large e^\frac{x}{y}=6x−6y\]????

OpenStudy (unknownrandom):

Yes

OpenStudy (irishboy123):

taken logs?

OpenStudy (unknownrandom):

\[e^{x/y}\frac{ y-x \frac{ dy }{ dx } }{ y^2 }=6-6\frac{ dy }{ dx}\]

OpenStudy (unknownrandom):

I have, but I don't remember them to well.

OpenStudy (irishboy123):

\(\Large f (x,y) = e^\frac{x}{y}- 6x+ 6y = 0\) \(\Large \nabla f = <\frac{1}{y} e^{\frac{x}{y}} - 6 , -\frac{x}{y^2} e^{x/y} +6>\) \[\Large y' = -\dfrac{f_x}{f_y} = \dfrac{y e^{\frac{x}{y}} - 6y^2}{x e^{\frac{x}{y}} +6y^2} \]

OpenStudy (irishboy123):

soz, correction: \[\Large \dfrac{dy}{dx} = -\dfrac{f_x}{f_y} = \color{red}{-}\dfrac{y e^{\frac{x}{y}} - 6y^2}{\color{red}{-}x e^{\frac{x}{y}} +6y^2} \\ \Large = \dfrac{6y^2 - y e^{\frac{x}{y}} }{6y^2-x e^{\frac{x}{y}} }\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

that is one way to do it you could do it without taking the log as well

OpenStudy (misty1212):

the derivative of the right side is easy, it is \(6-6y'\)

OpenStudy (misty1212):

the derivative of \[\frac{x}{y}\] is \[\frac{y-xy'}{y^2}\]

OpenStudy (misty1212):

putting it together gives \[\frac{y-xy'}{y^2}e^{\frac{x}{y}}=6-6y'\]

OpenStudy (unknownrandom):

@misty

OpenStudy (unknownrandom):

Can you break it down more?

OpenStudy (misty1212):

sure

OpenStudy (misty1212):

which part is not clear? it is clear that taking the derivative of \[6x-6y\] with respect to \(x\) gives \[6-y'\]?

OpenStudy (unknownrandom):

Yes. I am just having a hard time getting it to explicit form. y'=

OpenStudy (misty1212):

oh that that is a raft of algebra

OpenStudy (misty1212):

\[\frac{y-xy'}{y^2}e^{\frac{x}{y}}-6y'=6-6y'\] maybe start with \[\frac{y-xy'}{y^2}e^{\frac{x}{y}}-6y'=6\]

OpenStudy (misty1212):

break it up as \[\frac{y}{y^2}e^{\frac{x}{y}}-\frac{xyy'}{y^2}e^{\frac{x}{y}}-6y'=6\]

OpenStudy (misty1212):

subtract and get \[-\frac{xy'}{y}-6y'=6-\frac{1}{y}e^{\frac{x}{y}}\]

OpenStudy (misty1212):

factor out the \(y'\) on the left, then divide

OpenStudy (misty1212):

oh i made a mistake in the algebra

OpenStudy (misty1212):

\[-\frac{xy'}{y^2}-6y'=6-\frac{1}{y}e^{\frac{x}{y}}\]

OpenStudy (unknownrandom):

I got it. Thanks!